# A Systematic Literature Review on Vehicular Collaborative Perception – A Computer Vision Perspective

### v1

@@ VIII-A4 Simulation Experiment
In simulation experiments, trends similar to those observed in real-world evaluations are evident. Table XXII summarizes studies employing simulation-based experiments, which either generate new datasets or adapt existing ones to evaluate specific approaches, as demonstrated in [14].
@@
TABLE XXII: Overview of simulation experiments. Supported tasks are object detection (OD), object tracking (OT), semantic segmentation (SS), motion prediction (MP), lane detection (LD), map fusion (MF). A “-” indicates the absence of information in the referenced publication.
@@ VIII-B1 Environmental Settings
The methodologies for evaluating CP algorithms, including datasets and experiments, were introduced in the previous section. This section examines the specific scenarios used for algorithm evaluation. In both real-world and simulation studies, environments are typically categorized into three primary types: urban, rural, and highway, as outlined in Table XXIII. Simulation studies offer a wider range of scenarios compared to real-world evaluations, largely due to the extensive use of CARLA and its pre-defined maps. However, this reliance on CARLA maps introduces limitations in the diversity of road environments and the assessment of specific road features, such as intersections. Although many studies specify which CARLA map is utilized, detailed information about the types of intersections or the specific routes examined is often absent. Commonly evaluated road configurations, including cross intersections and straight road segments, are summarized in Table XXIV.
TABLE XXIII: Overview of the evaluation environments for real-world and simulation scenarios.
@@
TABLE XXIV: Overview of Road Environment for Evaluation
@@ VIII-C1 General Evaluation Metrics for Perception Tasks
The general evaluation metrics for different perception tasks are summarized in the Table XXV. These metrics are adapted from single-entity perception and are widely accepted by researchers. In some cases, evaluation results are divided into different groups based on detection difficulty levels, such as easy, medium, and difficult, as seen in the KITTI dataset [130], which considers factors like occlusion level and object size. Evaluation results can also be categorized by object type, such as cars, cyclists, and pedestrians. This categorization helps researchers better understand the strengths and limitations of different approaches.
TABLE XXV: Overview of general evaluation metrics for collaborative perception tasks
@@ VIII-C2 Custom Evaluation Metrics for Collaborative Perception
Traditional evaluation metrics used for single-entity perception do not adequately represent the performance of cooperative perception (CP). Since CP primarily aims to address visual occlusion problems and serves as a supplement to single-entity perception, it requires distinct evaluation criteria. Moreover, CP is significantly constrained by communication resources. Therefore, communication factors should be incorporated into the design of evaluation metrics. In this subsection, we summarize custom metrics designed for CP in Table XXVI. These metrics are classified into the following three types.
TABLE XXVI: Overview of custom evaluation metrics for collaborative perception tasks
@@ VIII-C3 Ablation Studies
TABLE XXVII: Ablation studies grouped by Aspect
@@ IX-C Risks
  • Deployment gap: A significant risk is the gap between research advancements and real-world deployment. Although various studies address individual aspects of this gap, the complexities of real-world conditions far exceed those modeled in datasets or simulations. For instance, unpredictable communication latency and interruptions can adversely affect CP performance. Additionally, striking the right balance between the perceptual improvements gained through collaboration and the additional costs incurred is challenging. Successful deployment of CP also requires collaboration among different vendors to ensure that CAVs from various manufacturers can communicate effectively and have compatible perception modules.
@@ Acknowledgments
TABLE XXVIII: Overview of the methods for collaborative object detection (COD) based on BEV and 3D representations. V: Vehicle, I: Infrastructure, Raw: Raw data fusion, Trad Feat: Traditional Feature Fusion, Atten Feat: Attention Feature Fusion, Obj Fusion: Object-level Fusion, Graph: Graph-based Fusion.
@@
TABLE XXIX: Overview of the methods for collaborative object detection (COD) based on BEV and 3D representations. (Continued)

### v3

@@ VIII-A4 Simulation Experiment
In simulation experiments, trends similar to those observed in real-world evaluations are evident. Table XXIII summarizes studies employing simulation-based experiments, which either generate new datasets or adapt existing ones to evaluate specific approaches, as demonstrated in [14].
@@
TABLE XXIII: Overview of simulation experiments. Supported tasks are object detection (OD), object tracking (OT), semantic segmentation ($S$), motion prediction (MP), lane detection (LD), map fusion (MF). A “-” indicates the absence of information in the referenced publication.
@@ VIII-B1 Environmental Settings
The methodologies for evaluating CP algorithms, including datasets and experiments, were introduced in the previous section. This section examines the specific scenarios used for algorithm evaluation. In both real-world and simulation studies, environments are typically categorized into three primary types: urban, rural, and highway, as outlined in Table XXIV. Simulation studies offer a wider range of scenarios compared to real-world evaluations, largely due to the extensive use of CARLA and its pre-defined maps. However, this reliance on CARLA maps introduces limitations in the diversity of road environments and the assessment of specific road features, such as intersections. Although many studies specify which CARLA map is utilized, detailed information about the types of intersections or the specific routes examined is often absent. Commonly evaluated road configurations, including cross intersections and straight road segments, are summarized in Table XXV.
TABLE XXIV: Overview of the evaluation environments for real-world and simulation scenarios.
@@
TABLE XXV: Overview of Road Environment for Evaluation
@@ VIII-C1 General Evaluation Metrics for Perception Tasks
The general evaluation metrics for different perception tasks are summarized in the Table XXVI. These metrics are adapted from single-entity perception and are widely accepted by researchers. In some cases, evaluation results are divided into different groups based on detection difficulty levels, such as easy, medium, and difficult, as seen in the KITTI dataset [133], which considers factors like occlusion level and object size. Evaluation results can also be categorized by object type, such as cars, cyclists, and pedestrians. This categorization helps researchers better understand the strengths and limitations of different approaches.
TABLE XXVI: Overview of general evaluation metrics for collaborative perception tasks
@@ VIII-C2 Custom Evaluation Metrics for Collaborative Perception
Traditional evaluation metrics used for single-entity perception do not adequately represent the performance of cooperative perception (CP). Since CP primarily aims to address visual occlusion problems and serves as a supplement to single-entity perception, it requires distinct evaluation criteria. Moreover, CP is significantly constrained by communication resources. Therefore, communication factors should be incorporated into the design of evaluation metrics. In this subsection, we summarize custom metrics designed for CP in Table XXVII. These metrics are classified into the following three types.
TABLE XXVII: Overview of custom evaluation metrics for collaborative perception tasks
@@ VIII-C3 Ablation Studies
TABLE XXVIII: Ablation studies grouped by Aspect
@@ IX-C Risks
  • Deployment gap: A significant risk is the gap between research advancements and real-world deployment. Although various studies address individual aspects of this gap, the complexities of real-world conditions far exceed those modeled in datasets or simulations. For instance, limited communication bandwidth caused by environmental factors, unexpected synchronization failures, as well as unpredictable communication latency and interruptions can adversely affect CP performance. Additionally, striking the right balance between the perceptual improvements gained through collaboration and the additional costs incurred is challenging. Successful deployment of CP also requires collaboration among different vendors to ensure that CAVs from various manufacturers can communicate effectively and have compatible perception modules.
@@ Acknowledgments
TABLE XXIX: Overview of the methods for collaborative object detection (COD) based on BEV and 3D representations. V: Vehicle, I: Infrastructure, Raw: Raw data fusion, Trad Feat: Traditional Feature Fusion, Atten Feat: Attention Feature Fusion, Obj Fusion: Object-level Fusion, Graph: Graph-based Fusion.
@@
TABLE XXX: Overview of the methods for collaborative object detection (COD) based on BEV and 3D representations. (Continued)